# Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v4

@@ 2.3.2 Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected the five best-performing prompts. Fig. 2 illustrates the construction strategies for the five prompts.
  • Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
@@
  • Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$). We acknowledge that providing a list of CWEs may limit the scope of detection set of CWEs and miss security defects that do not fall into the provided CWE categories. However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
  • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
@@
Table 2: Adjusted IHMU-category for LLM responses.
Classification Definition Instance
Instrumental (I) The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario. Reviewer: (Line 76)I think this is race-y. Let’s say we have a three coroutines scheduled to call function. First one gets to L74 and function yields. Second coroutine gets in before threshold and gets to L68, sets is_deferred and sleeps on L70. First coroutine is finished and unsets on L76. Third coroutine enters still before threshold so L62 doesn’t influence it as planned and gets to L70, sleeps similar time as second coroutine. Then both, second and third coroutines call function in very short time difference.
LLM: **Potential Time of Check to Time of Use (TOCTOU) race condition in ‘Throttler.call()‘** (Lines 58-77):
@@
- **Solution**: A synchronization mechanism (like a lock) should be employed to ensure thread safety.
Explanation: LLM indicated the security defect identified by the reviewer, with the correct type, accurate code line range and detailed description, thus categorized as Instrumental.
Helpful (H) The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate. However, the response still provides useful evidence pointing toward the existence of the target defect. Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g. password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is. But I guess that’s unlikely.
LLM:**Potential Information Disclosure**(Line 141):
@@
- **Solution**: Limit the details being logged and ensure that all sensitive data is properly obfuscated before logging.
Explanation: The reviewer indicated that the mask_password function might incorrectly replace different types of sensitive information. While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario. Consequently, this instance was categorized as Helpful rather than Instrumental.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task. Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole. In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk. This would result in copying the compute host’s root disk to the specified rbd volume and exposing it to the user.
@@
Explanation: LLM did not detect any security defects, but the reviewer actually identified a security hole in the disk format detection. Therefore, this instance was considered Misleading.
Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed. Reviewer: (Line 469) So not even timeout cancels the future? Will this cause that threads will hang around forever leaking OS resources? This feels dangerous.
@@
2.3.3 Evaluation Metrics
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al. (2020) for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al. (2020) because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain) detailed in Table 2. To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies fewer misleading results produced during security defect detection. Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
@@ 2.3.4 Baselines
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Marjamäki et al., 2024) for C/C++ and Bandit (PyCQA, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (Dennison et al., 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
@@ 2.3.5 Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ datasets. We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 3.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
@@ 2.3.6 Consistency Calculation
To measure the non-determinism of LLMs, we adopted the metrics in Chang et al. (2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
@@ 2.4 Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses. We employed the open coding and constant comparison (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. In particular, we excluded the Consistency theme defined in Kabir et al. (2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, ten responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the two authors completed the remaining data analysis and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset. To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by Huang et al. (2025), which categorizes hallucinations into two main types: Factual and Faithfulness. Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication. Faithfulness hallucination, on the other hand, refers to deviations of the output from the instruction, current context, and reasoning process, including Instruction Inconsistency, Context Inconsistency, and Logical Inconsistency. We mapped the quality problems defined in our codebook to these hallucination subtypes to calculate their distribution across the dataset. Detailed mapping and results are provided in Section  3.2.
@@ 2.5 Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
@@
Token Adjusted from  Paul et al. (2021); McIntosh et al. (2016) Number of tokens of the code file. Computed using the TikToken OpenAI (2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
@@
SecurityDefectType Transformed from  Chen et al. (2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. Yu et al. (2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities Chen et al. (2023); Zhou et al. (2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  Paul et al. (2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  Paul et al. (2021); McIntosh et al. (2016) McCabe’s Cyclomatic Complexity McCabe (1976). Computed using the Lizard Terry (2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases Paul et al. (2021).
@@
AuthorExperience Collected from  Paul et al. (2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
AnnotationHasCode Transformed from  Chen et al. (2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs Chen et al. (2023), leading to incorrect security defect detection.

### v5

@@ 2.3.2 Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (liu2023pre). To design prompts, we followed the best practices published by OpenAI (GPT2023bestpractices) and the three prompt design strategies in zhang2024prompt: basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected the five best-performing prompts. Fig. 5 illustrates the construction strategies for the five prompts.
  • Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.’. It is worth noting that we instruct the model to generate a solution for each identified security defect, since an incorrectly detected security defect usually cannot lead to a logically consistent fix in the source code. This process naturally drives the model to evaluate the validity of its detection and acts as an implicit self-reflection mechanism.
@@
  • Prompt4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (pan2023fine)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$). We acknowledge that providing a list of CWEs may limit the scope of detection set of CWEs and miss security defects that do not fall into the provided CWE categories. However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
  • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$$\textbf{P}_{cot-2}$. Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
However, in the two-stage $\textbf{P}_{cot}$, the performance of LLM can be influenced by three factors: the commit message, CoT guidance, and the code context generated by LLM. Among these, the fabricated code context may contradict the actual code, leading to incorrect judgments. To isolate the impact of the LLM-generated context, we proposed a guardrail version of the $\textbf{P}_{cot}$ prompt and conducted experiments exclusively on DeepSeek-R1 under this additional prompt.
  • Prompt 6 ($\textbf{P}_{cot-guardrail}$): This prompt retains the commit message and CoT instruction but does not use the code context generated by LLM, allowing us to evaluate the effect of fabricated contexts. The detailed template is shown in Fig 5.
@@
Figure 5: Construction template for the guardrail version of $\textbf{P}_{cot}$
2.3.3 Evaluation Metrics
During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements. We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by mahajan2020recommending for measuring responses in Stack Overflow. We adapted the evaluation metrics formulated in their work (mahajan2020recommending) as follows: First, we excluded the ‘Unavailable’ category defined by mahajan2020recommending because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain). To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description. The detailed definition and corresponding examples of these four rating categories are as follows:
Instrumental (I): The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario.
Example 1 - Instrumental Reviewer: (Line 76)I think this is race-y. Let’s say we have a three coroutines scheduled to call function. the first may yield at L74, the second enters before the threshold and sleeps at L70 while setting is_deferred, then the first finishes and unsets at L76, allowing the third to enter before the threshold as well. Then both, second and third coroutines call function in very short time difference.
LLM: Race condition in ‘Throttler.call()‘(Lines 58-77):
@@ 2.3.3 Evaluation Metrics
Explanation: LLM indicated the race condition defect identified by the reviewer, with the accurate code line range and detailed description, thus categorized as Instrumental.
Helpful (H): The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate. However, the response still provides useful evidence pointing toward the existence of the target defect.
Example 2 - Helpful Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g. password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is. But I guess that’s unlikely.
LLM:Potential Information Disclosure(Line 141):
@@
Explanation: The reviewer noted that the mask_password function might incorrectly replace different types of sensitive information. While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario. So this instance was categorized as Helpful rather than Instrumental.
Misleading (M): The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
Example 3 - Misleading Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole. In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk. This would result in copying the compute host’s root disk to the specified rbd volume and exposing it to the user.
@@
Explanation: The reviewer pointed out a sensitive data exposure at line 329, whereas the LLM did not detect any security defects. Therefore, it was classified as Misleading.
Uncertain (U): The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
Example 4 - Uncertain Reviewer: (Line 469) So not even timeout cancels the future? Will this cause that threads will hang around forever leaking OS resources? This feels dangerous.
@@
The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers. At the same time, a lower M-Score signifies fewer misleading results produced during security defect detection. Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
It is important to note that the cost and latency of LLMs in security code review are significant concerns for practitioners. However, in this study, these two points were not specifically discussed. Due to the various invocation methods of the LLMs involved (including calls through the ChatGPT platform, local deployment, and API calls), it is difficult to provide a consistent cost estimate across different LLMs. Moreover, the latency of the models is significantly affected by their deployment environments, such as computational resources, inference engines, and parallelization strategies, making the latency of different LLMs in our experiments incomparable.
@@ 2.3.4 Baselines
We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (ren2024prorlearn), GRACE (lu2024grace) and LineVul (fu2022linevul). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (cppcheck) for C/C++ and Bandit (bandit) for Python. Furthermore, we also used CodeQL (codeql) and SonarQube (sonarqube) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (semgrep) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
@@ 2.3.5 Data Labelling
[…]
@@ 2.3.6 Consistency Calculation
To measure the non-determinism of LLMs, we adopted the metrics in chang2024beyond and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, an LLM generated three responses for a given code file under a specific prompt template. We then calculated the entropy for the four response categories in the three responses. We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher mean entropy suggests the LLM has higher inconsistency in repetitions of experiments. To further quantify the uncertainty of the entropy estimates and ensure the reliability of our findings, bootstrap confidence intervals were calculated for each LLM–prompt combination. Following the guidance of  efron1994introduction, we resampled the original dataset with replacement to match its original size and calculated the mean entropy for each resample. As suggested by  davison1997bootstrap, if confidence levels 0.95 and 0.99 are to be used, then it is advisable to have [the number of resamples]=999 or more, if practically feasible. Therefore, consistent with the settings adopted in  ahmed2025can, we performed 1,000 bootstrap replications to obtain the 95% confidence intervals for the entropy estimates.
@@ 2.4 Research Procedure of RQ2
[…]
We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset. To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by huang2025survey, which categorizes hallucinations into two main types: Factual and Faithfulness. Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication. Faithfulness hallucination, on the other hand, refers to deviations of the output from the instruction, current context, and reasoning process, including Instruction Inconsistency, Context Inconsistency, and Logical Inconsistency. We mapped the quality problems defined in our codebook to these hallucination subtypes to calculate their distribution across the dataset. Detailed mapping and results are provided in Section  3.2.
@@ 2.5 Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (harrell2001regression) to construct a regression model. As mentioned in Section 2.3.3, the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (powers2008statistical). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is described in the following subsections.
@@
Token Adjusted from  paul2021security; mcintosh2016empirical Number of tokens of the code file. Computed using the TikToken tiktoken tokenizer. Larger code files may be more difficult for LLMs to analyze.
Position Collected from  Sovrano_2025 Position of the security defect within the code file. Determined by calculating the proportion of tokens preceding the defect relative to the total token count, using the TikToken tiktoken tokenizer. LLMs tend to significantly underperform when detecting security defects located toward the end of files (Sovrano_2025; rafi2024impact).
@@
SecurityDefectType Transformed from  chen2023chatgpt Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by yu2023security. ChatGPT is proven to display distinct detection performances across various vulnerabilities chen2023chatgpt; zhout2023devil. Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit Collected from  paul2021security Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity Collected from  paul2021security; mcintosh2016empirical McCabe’s Cyclomatic Complexity mccabe1976complexity. Computed using the Lizard lizard analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases paul2021security.
@@
AuthorExperience Collected from  paul2021security Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
AnnotationHasCode Transformed from  chen2023chatgpt Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs chen2023chatgpt, leading to incorrect security defect detection.